Two point charges Q1 = +3.2 ?C and Q2 = +3.6 ?C are initially separated by 1.0 m and held fixed. The charges are then released, and both move in response to the electric force between them. What is the sum of the kinetic energies of the two particles when they are very far apart?
Answer #1

To know the sum of kinetic energy we use the conservation of energy

at the initial position we have:

E_0 = \frac{kQ_1Q_2}{d} = \frac{9\times 10^9(3.2\times 10^{-6})(3.6\times 10^{-6})}{1} = 0.10368 J

Whern they are far away:

E_0 = K_1 + K_2

so:

E_0 = E_f

K_1 + K_2 = 0.10368 J